OpenStudy (anonymous):

For Introduction to Statistics and Probability, how do you a sample variance and sample population?

OpenStudy (anonymous):

Sample variance is: \[s^2 = \frac{1}{n - 1} \sum_{i=1}^n (x_i - m)^2\] And there really isn't a "sample population". A sample is taken from a population, since you cannot test and entire population. You collect just a sample size in lieu of testing the whole population.

OpenStudy (anonymous):

Thank you!